# Excitation Amplitude Sampling for Low Variance Electronic Structure on Quantum Computers

### v1

@@ III Non-energetic properties
Figure 3: Systematic and random errors in the magnetic correlations of Hydrogen chains (a) The spin-spin correlation function of an eight atom hydrogen chain at $1.5$Å  symmetric stretching between the first atom and its $i^{\text{th}}$ nearest neighbor in a STO-3G basis. Hartree–Fock (light blue), FCI (red), and CCSD (grey) are shown as classical methods, while ten sets of 40,960 shots are used to estimate the mean and error of the expectation values from the exact statevector, comparing a direct sampling of qubitwise grouped Pauli strings (purple) and a mapping of the shadow state to a CCSD ansatz (orange). (b) The local (same-atom) spin-spin correlation function for an Natom hydrogen chain, showing the standard deviation of the distribution of expectation values with 40,960 shots for the two approaches, with the navy dashed line also showing the systematic error in the CCSD model for this property when the exact coefficients are used, quantifying the error introduced by this CCSD model parameterization.
@@
This allows us to use the previously obtained excitation amplitudes from the shadow state in order to define this classically tractable coupled cluster surrogate model, from which properties of interest can be computed efficiently. Importantly, this coupled cluster $|\Psi_{\text{cc}}\rangle$ state will share the same energy as the state $|\Psi\rangle$, which will therefore be exact if the exact excitation amplitudes $c_{i}^{a}$ and $c_{ij}^{ab}$ are used for this mapping. This property is exploited in ‘tailored’ coupled cluster approaches, which have recently been combined with quantum algorithms in order to introduce quantum-derived information within a low-energy subspace into a classical CCSD state [76, 77]. However, in this work we are defining a coupled cluster state parameterization over the entire space as a surrogate model of the quantum state, and considering the accuracy to which properties can be estimated from this tractable form.
@@
$\Gamma_{\alpha\beta\gamma\delta}=\braket{\Psi}{\hat{c}^{\dagger}_{\alpha}\hat{c}^{\dagger}_{\beta}\hat{c}_{\delta}\hat{c}_{\gamma}}{\Psi}.$ (23)
@@
$\Gamma_{\alpha\beta\gamma\delta}=\braket{\phi}{(1+\hat{\Lambda})e^{-\hat{T}}\hat{c}^{\dagger}_{\alpha}\hat{c}^{\dagger}_{\beta}\hat{c}_{\delta}\hat{c}_{\gamma}e^{\hat{T}}}{\phi},$ (24)
@@
Due to the truncation of the cluster operator to only single and double excitations in this work (Eq. 22), the properties derived from this classical surrogate state will incur a systematic error from the incomplete state definition, even if the excitation amplitudes from which the state is constructed are themselves exact. This contrasts with the properties of the mixed energy estimator. Nevertheless, the accuracy by which the properties can be extracted in principle surpasses that of a classical optimization of the coupled-cluster state, since the amplitudes are appropriately relaxed due to the correlations which manifest beyond the truncated excitation level considered. Furthermore, it is possible to systematically alleviate the constraints of the model by increasing rank of excitations which are considered in the $\hat{T}$ operator of Eq. 22, which would become exact (and therefore the properties equivalent to those of the Pauli estimator) in the case that all excitations are considered. However, we restrict to a single and double excitation parameterization of the coupled-cluster model in this work.
To numerically test the performance of these non-energetic properties from this CCSD surrogate model, we return to the chain of eight Hydrogen atoms and consider the magnetic spin correlation functions which emerge from the correlated electronic structure upon symmetrically stretching the atoms to $1.5$Å. This question has been heavily studied by classical electronic structure methods to characterize long-range magnetic order in this system [104], and is a central observable in many (e.g. neutron scattering) experimental setups. We define the two-point spin correlation function between two Hydrogen atoms, $H_{i}$ and $H_{j}$, as $\braket{\hat{S}_{z}(H_{i})\hat{S}_{z}(H_{j})}$, with the spin magnetization operator for the atom $H_{i}$ defined as
@@
In Fig. 3(a) we calculate this spin-spin correlation function between the first atom and increasingly distant atoms in the chain. The plot shows the results of a Hartree-Fock calculation (light blue), which largely misses the correlated spin fluctuations of the electrons between atoms, or any long-range antiferromagnetic (AFM) order which is shown to emerge in the exact FCI (red) result. We also show the results of a classically optimized CCSD state (grey), which despite being accurate for on-atom and nearest-neighbor magnetic fluctuations, does not accurately describe the decay of the correlations with distance, exhibiting an erroneously slow decay of this AFM order. After encoding the exact state on a statevector simulator, we can also compare the approaches to sampling this observable as would be done on a quantum device. We can evaluate this spin correlator by sampling the corresponding Pauli operators directly, grouped qubitwise into mutually commuting sets (purple), or by sampling excitation amplitudes from the classical shadow procedure and mapping to the classical CCSD surrogate state as described above (orange). Both used a total of ten sets of 40,960 shots to estimate the error, allowing a faithful comparison between these statistically sampled approaches.
@@
As expected, directly measuring the Pauli strings of the correlation function incurs no systematic error, with the exact results within statistical errors over all length scales. This contrasts with the approach relying on mapping the sampled state to a CCSD model, which exhibits a systematic error originating from the incompleteness of the ansatz, as well as covariances arising from products of stochastic $t$-amplitudes in the non-linear functional for the RDM of Eq. 24. However, we find that this systematic error is always less than the classically optimized CCSD state, and results in a reduction in the particularly poor description of the long-range correlations by a factor of two, partially correcting the deficiency in the classical CCSD description. Furthermore, we find a large reduction in the statistical errors associated with this estimate compared to the Pauli estimator.
@@
Whether this approach and the level of systematic error is acceptable compared to the Pauli estimator will be dependent on the system, accuracy required, and the level of stochastic error in the Pauli estimator for a given shot budget. To consider this, and the scaling of the results with system size, in Fig. 3(b) we consider the same-atom spin-spin correlations, $\langle\hat{S}_{z}(H_{0})\hat{S}_{z}(H_{0})\rangle$, and decompose the error from the CCSD surrogate model into its systematic and random components, and compare these to the random error of the Pauli estimator for this local expectation value for the same number of shots. These errors are then compared as the system grows in size as more atoms are included for a fixed interatomic distance. For this correlation function the statistical error in the Pauli estimator grows significantly quicker over the accessible system sizes compared to both the statistical error and systematic error of using the CI amplitudes measured on the quantum device and mapped to the CCSD ansatz. In this case therefore, the systematic error is small enough that it is well inside the larger random errors of the Pauli estimators, clearly demonstrating a benefit in the approach. Furthermore, both the random and systematic errors in the mapped CCSD state appear to grow linearly, compared to the expected higher power of system size for the Pauli estimator. However, we expect the effectiveness of this approach to be dependent on the correlation function of interest, with Fig. 3(a) already demonstrating that the larger distance correlation functions exhibit a larger systematic error. Nevertheless, a systematic improvement in this error can always be found via the inclusion of higher order excitations in the coupled cluster model, which will be considered in the future.
@@ IV Application to quantum embedding
We can also go beyond exact statevector cluster solutions, and consider a full NISQ workflow using VQE as the cluster solver, followed by the shadow tomography of the excitation amplitudes of each cluster and recombining for the full system energy. In Fig. 5 we apply this to cubic Boron Nitride in an FCC lattice (space group $F{\overline{4}}3m$), considering atomic fragments consisting of the $2p$ orbitals on each atom, giving a cluster size of 12 spin-orbitals, and the core $(1s,2s)$ orbitals of each atom in the primitive cell treated separately. Within a $3\times 3\times 3$ $k$-point sampled cell, each cluster was solved with VQE with the unitary coupled-cluster (UCCSD) ansatz. Only one optimization per cluster was required, with 4,000 shots used to build a classical shadow and extract out the full system mixed energy estimator. Twelve independent samples of the shadow state were taken to obtain a mean and variance on the energy estimate, with error bars barely visible on the plot. The equivalent FCI cluster results are always within two standard deviation and the error can be systematically reduced with more samples and improved ground state preparation algorithms. We find that the effect of the correlated treatment is to enlarge the lattice parameter to coincide with the experimental result of $c=3.615$Å [131]. The correlations in this system are relatively modest, yet the reliability and robustness of the overall protocol remains, allowing smooth equations of state even when combined with approximate state preparation quantum algorithms. This allows us to go beyond the largely single-geometry energy estimates of previous approaches in combining embedding approaches with quantum solvers.
@@
Acknowledgements
The authors wish to thank Mohammadreza Mousavi for helpful comments over the course of this work. This research has been partially supported by the EPSRC project on Verified Simulation for Large Quantum Systems (VSL-Q), grant reference EP/Y005244/1 and the EPSRC project on Robust and Reliable Quantum Computing (RoaRQ), Investigation 009 Model-based monitoring and calibration of quantum computations (ModeMCQ), grant reference EP/W032635/1. We acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team.
@@ Appendix B Hamiltonian dependent variance factor
Figure 6: Hamiltonian dependent scaling of the variance prefactors. The Hamiltonian dependent factor in the variance expression for the Pauli (blue) and mixed (orange) estimators for a symmetrically stretched STO-3G basis Hydrogen chain with $\AA$ interatomic distance.
@@
where the variance of the Pauli string, $\hat{P}_{a}$, is bounded $10<\mathrm{Var}[\hat{P}_{a}]<$ 1. For the mixed correlation energy estimator, the scaling in shot number instead depends on the difference of electron repulsion integrals, $v_{ijab}$, as
@@
To numerically compare these quantities and investigate how they depend on system size, in Fig. 6 we plot a comparison of the terms in Eqs.29 and 30 for increasing length of hydrogen chains in a minimal basis with an atomic spacing of $1.5$Å in a canonical basis representation. The variance of each Pauli string in Eq. 29 is unknown and dependent on the state of interest. We therefore assume that the $\mathrm{Var}[\hat{P}_{a}]=1$ for all Pauli strings, and take the weights of the Hamiltonian, $w_{a}$, to arise from the Jordan-Wigner fermion-to-qubit mapping. With this simplifying assumption, the prefactors are both found to increase linearly with system size, albeit with a substantially smaller gradient for the mixed estimator prefactor. This validates the arguments of the main text that the correlation energy estimator has a smaller prefactor, encoding just the electron repulsion terms rather than all the physics of the Hamiltonian. We also note this is just the prefactor scaling, with the full scaling in shot number for the Pauli estimator then multiplied by a polynomial factor, while for the mixed estimator the explicit system size dependent factor is logarithmic, combining to result in the favorable scaling observed in Fig. 1(b).
Appendix C Calculation of stabilizer overlaps
@@
In this work, we are restricted to considering |$\ket{\phi_{i}}$ (as well as |$\ket{b_{j}}$) to be computational basis states, with $U_{j}$ an $N$-qubit Clifford unitary, amounting to the calculation of the overlap (including its phase) between a stabilizer state and a computational basis state. An approach for this calculation is given in Ref. [134], however, given the simplification of only requiring overlap with a computational basis state we can simplify the algorithm to avoid tracking of the global phase of the stabilizer state.
To perform this calculation we first reduce the stabilizer matrix to a canonical form [134]. The absolute value of the inner product of the stabilizer state with each computational basis state is given by $2^{-s/2}$ where $s$ is the $X$-rank of the reduced stabilizer matrix (the number of generators containing an X or Y Pauli) [134]. To calculate the phase we then perform a measurement of the circuit which creates the stabilizer state $U_{j}^{\dagger}\ket{b_{j}}$ using the algorithm shown in ref. [97]. This results in one of the computational basis states which overlap with the stabilizer state $U_{j}^{\dagger}\ket{b_{j}}$. From this measured state we can use the generators in the canonical stabilizer matrix to transform it into each of the basis states of interest, i.e. |$\ket{0}$ and |$\ket{\phi_{i}}$, while tracking the phase generated by the action of each generator on the state. The Gaussian elimination step scales as $\mathcal{O}(N^{3})$, which is the most costly step of the algorithm and it must be repeated for each measurement in the shadow $U_{i}^{\dagger}\ket{b_{i}}\bra{b_{i}}U_{i}$. The overall cost of the post processing will therefore be determined by the number of basis states that the phase needs to be calculated for, resulting in an overall scaling in order to compute the singles and doubles excitation amplitudes of $\mathcal{O}(M(N^{4}+N^{3}))$.
@@
Appendix D Direct configurational sampling
Figure 7: Variance scaling for mixed and Pauli estimators of Hydrogen chains over changing correlation strengths and chain length, analogous to Fig. 1, but including a direct sampling of excitation amplitudes in the mixed estimator. (a) Comparison of the distributions of the Pauli and mixed energy estimators, comparing classical shadows to the direct evaluation of the excitation amplitudes via diagonalization and measurement of the operators in Eq. 32 (green), for the ten atom STO-3G basis Hydrogen chain across a range of interatomic distances. The width of the error bar is the standard deviation of the distribution of energies from $1000$ measurements of the exact statevector. (b) The dependence of the standard deviation of each energy estimator with system size, for a fixed atomic distance of $1.5$Åand $1000$ shots. The black line again shows the Pauli operators grouped into as few commuting sets as possible, while purple is the same for grouping Pauli operators based on qubit wise commutivity. The orange and green lines represent the standards deviations of the mixed estimator using the direct and Clifford shadows methods of calculating the $c_{ij}^{ab}$ amplitudes.

### v2

@@ III Non-energetic properties
Figure 3: Systematic and random errors in the magnetic correlations of Hydrogen chains (a) The spin-spin correlation function of an eight atom hydrogen chain (symmetrically stretched to 1.51.5Å) between the first atom and its $i^{\text{th}}$ nearest neighbor in a STO-3G basis. Hartree–Fock (light blue), FCI (red), and CCSD (grey) are shown as classical methods, while ten sets of 40,960 shots are used to estimate the mean and error of the expectation values from the exact statevector, comparing a direct sampling of qubitwise grouped Pauli strings (purple) and a mapping of the shadow state to a CCSD ansatz (orange). (b) The local (same-atom) spin-spin correlation function for an Natom hydrogen chain, showing the standard deviation of the distribution of expectation values with 40,960 shots for the two approaches, with the navy dashed line also showing the systematic error in the CCSD model for this property when the exact coefficients are used, quantifying the bias introduced by this CCSD model parameterization.
@@
This allows us to use the previously obtained excitation amplitudes from the shadow state in order to define this classically tractable coupled cluster surrogate model, from which properties of interest can be computed efficiently. Importantly, this coupled cluster $|\Psi_{\text{cc}}\rangle$ state will share the same energy as the state $|\Psi\rangle$, which will therefore be exact if the exact excitation amplitudes $c_{i}^{a}$ and $c_{ij}^{ab}$ are used for this mapping. This property is exploited in ‘tailored’ coupled cluster approaches, which have recently been combined with quantum algorithms in order to introduce quantum-derived information within a low-energy subspace into a classical CCSD state [76, 77]. However, in this work we are defining a coupled cluster state parameterization over the entire space as a surrogate model of the quantum state, and considering the accuracy to which properties can be estimated from this classically tractable form.
@@
$\Gamma_{\alpha\beta\gamma\delta}=\braket{\Psi|\hat{c}^{\dagger}_{\alpha}\hat{c}^{\dagger}_{\beta}\hat{c}_{\delta}\hat{c}_{\gamma}|\Psi}.$ (23)
@@
$\Gamma_{\alpha\beta\gamma\delta}=\braket{\phi|(1+\hat{\Lambda})e^{-\hat{T}}\hat{c}^{\dagger}_{\alpha}\hat{c}^{\dagger}_{\beta}\hat{c}_{\delta}\hat{c}_{\gamma}e^{\hat{T}}|\phi},$ (24)
@@
Due to the truncation of the cluster operator to only single and double excitations in this work (Eq. 22), the properties derived from this classical surrogate state will incur a systematic error from the incomplete state definition, even if the excitation amplitudes from which the state is constructed are themselves exact. This contrasts with the properties of the mixed energy estimator. Nevertheless, the accuracy by which the properties can be extracted in principle surpasses that of a classical optimization of the coupled-cluster state, since the amplitudes are appropriately relaxed due to the correlations which manifest beyond the truncated excitation level considered. Furthermore, it is possible to systematically alleviate the constraints of the model by increasing rank of excitations which are considered in the $\hat{T}$ operator of Eq. 22, which would become unbiased in the case that all excitations are considered. However, we restrict to a single and double excitation parameterization of the coupled-cluster model in this work.
To numerically test the performance of these non-energetic properties from this CCSD surrogate model, we return to the chain of eight Hydrogen atoms and consider the magnetic spin correlation functions which emerge from the correlated electronic structure upon symmetrically stretching the atoms to 1.51.5Å. This question has been heavily studied by classical electronic structure methods to characterize long-range magnetic order in this system [109], and is a central observable in many (e.g. neutron scattering) experimental setups. We define the two-point spin correlation function between two Hydrogen atoms, $H_{i}$ and $H_{j}$, as $\braket{\hat{S}_{z}(H_{i})\hat{S}_{z}(H_{j})}$, with the spin magnetization operator for the atom $H_{i}$ defined as
@@
In Fig. 3(a) we calculate this spin-spin correlation function between the first atom and increasingly distant atoms in the chain. The plot shows the results of a Hartree-Fock calculation (light blue), which largely misses the correlated spin fluctuations of the electrons between atoms, or any long-range antiferromagnetic (AFM) order which is shown to emerge in the exact FCI (red) result. We also show the results of a classically optimized CCSD state (grey), which despite being accurate for on-atom and nearest-neighbor magnetic fluctuations, does not accurately describe the decay of the correlations with distance, exhibiting an erroneously slow decay of this AFM order. After encoding the exact state on a statevector simulator, we can also compare the approaches to sampling this observable as would be done on a quantum device. We can evaluate this spin correlator by sampling the corresponding Pauli operators directly, grouped qubitwise into mutually commuting sets (purple), or by sampling excitation amplitudes from the classical shadow procedure and mapping to the classical CCSD surrogate state as described above (orange). Both used a total of ten sets of 40,960 shots to estimate the bias from the CCSD mapping, as well as allowing a faithful comparison of the statistical errors between these two sampling approaches.
@@
As expected, directly measuring the Pauli strings of the correlation function incurs no systematic error, with the exact results within statistical errors over all length scales. This contrasts with the approach relying on mapping the shadow sampled state to a CCSD model, which exhibits a systematic bias originating from the incompleteness of the ansatz, as well as covariances arising from products of stochastic $t$-amplitudes in the non-linear functional for the RDM of Eq. 24. However, we find that this systematic error is always less than the classically optimized CCSD state, and results in a reduction in the particularly poor description of the long-range correlations by a factor of two, partially correcting the deficiency in the classical CCSD description. Furthermore, we find a large reduction in the statistical errors associated with this estimate compared to the Pauli estimator.
@@
Whether this approach and the level of systematic error is acceptable compared to the Pauli estimator will be dependent on the system, accuracy required, and the level of stochastic error in the Pauli estimator for a given shot budget. To consider this, and the scaling of the results with system size, in Fig. 3(b) we consider the same-atom spin-spin correlations, $\langle\hat{S}_{z}(H_{0})\hat{S}_{z}(H_{0})\rangle$, and decompose the overall error from the CCSD surrogate model into its systematic bias and variance, and compare these to the random error of the Pauli estimator for this local expectation value for the same number of shots. These errors are then compared as the system grows in size as more atoms are included for a fixed interatomic distance. For this correlation function the statistical error in the Pauli estimator grows significantly quicker over the accessible system sizes compared to both the statistical error and systematic error of using the CI amplitudes measured on the quantum device and mapped to the CCSD ansatz. In this case therefore, the systematic error is small enough that it is well inside the larger random errors of the Pauli estimators, clearly demonstrating a benefit in the approach. Furthermore, both the random and systematic errors in the mapped CCSD state appear to grow linearly, compared to the expected higher power of system size for the Pauli estimator.
However, we expect the effectiveness of this approach to be dependent on the correlation function of interest, with Fig. 3(a) already demonstrating that the larger distance correlation functions exhibit a larger systematic error. Nevertheless, a systematic improvement in this error can always be found via the inclusion of higher order excitations in the coupled cluster model, which will be considered in the future. For further comparison between these approaches, we can combine the bias and variance errors of the estimates into a mean squared error in the spin correlation functions for a given shot budget, which we provide in Appendix C for further insights into the efficiency of the CCSD mapped shadow approach.
@@ IV Application to quantum embedding
We can also go beyond exact statevector cluster solutions, and consider a full NISQ workflow using VQE as the cluster solver, followed by the shadow tomography of the excitation amplitudes of each cluster and recombining for the full system energy. In Fig. 5 we apply this to cubic Boron Nitride in an FCC lattice (space group $F{\overline{4}}3m$), considering atomic fragments consisting of the $2p$ orbitals on each atom, giving a cluster size of 12 spin-orbitals, and the core $(1s,2s)$ orbitals of each atom in the primitive cell treated separately. Within a $3\times 3\times 3$ $k$-point sampled cell, each cluster was solved with VQE with the unitary coupled-cluster (UCCSD) ansatz. Only one optimization per cluster was required, with 4,000 shots used to build a classical shadow and extract out the full system mixed energy estimator. Twelve independent samples of the shadow state were taken to obtain a mean and variance on the energy estimate, with error bars barely visible on the plot. The equivalent FCI cluster results are always within two standard deviation and the error can be systematically reduced with more samples and improved ground state preparation algorithms. We find that the effect of the correlated treatment is to enlarge the lattice parameter to coincide with the experimental result of c=3.615c=3.615Å [136]. The correlations in this system are relatively modest, yet the reliability and robustness of the overall protocol remains, allowing smooth equations of state even when combined with approximate state preparation quantum algorithms. This allows us to go beyond the largely single-geometry energy estimates of previous approaches in combining embedding approaches with quantum solvers.
@@
Acknowledgments
The authors wish to thank Mohammadreza Mousavi for helpful comments over the course of this work. This research has been partially supported by the EPSRC project on Verified Simulation for Large Quantum Systems (VSL-Q), grant reference EP/Y005244/1 and the EPSRC project on Robust and Reliable Quantum Computing (RoaRQ), Investigation 009 Model-based monitoring and calibration of quantum computations (ModeMCQ), grant reference EP/W032635/1. For the purpose of open access, the author has applied a Creative Commons Attribution (CC BY) license to any Author Accepted Manuscript version arising. The data that support the findings of this article are openly available [139]. We acknowledge the use of IBM Quantum services for this work. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team.
@@ Appendix B Hamiltonian dependent variance factor
[…]
@@
where the variance of the Pauli string, $\hat{P}_{a}$, is bounded $0<\mathrm{Var}[\hat{P}_{a}]<1$. For the mixed correlation energy estimator, the scaling in shot number instead depends on the difference of electron repulsion integrals, $v_{ijab}$, as
@@
To numerically compare these quantities and investigate how they depend on system size, in Fig. 6 we plot a comparison of the terms in Eqs.29 and 30 for increasing length of hydrogen chains in a minimal basis with an atomic spacing of 1.51.5Å in a canonical basis representation. The variance of each Pauli string in Eq. 29 is unknown and dependent on the state of interest. We therefore assume that the $\mathrm{Var}[\hat{P}_{a}]=1$ for all Pauli strings, and take the weights of the Hamiltonian, $w_{a}$, to arise from the Jordan-Wigner fermion-to-qubit mapping. With this simplifying assumption, the prefactors are both found to increase linearly with system size, albeit with a substantially smaller gradient for the mixed estimator prefactor. This validates the arguments of the main text that the correlation energy estimator has a smaller prefactor, encoding just the electron repulsion terms rather than all the physics of the Hamiltonian. We also note this is just the prefactor scaling, with the full scaling in shot number for the Pauli estimator then multiplied by a polynomial factor, while for the mixed estimator the explicit system size dependent factor is logarithmic, combining to result in the favorable scaling observed in Fig. 1(b).
Appendix C Mean squared errors of magnetic estimates of Hydrogen chains
Figure 7: Mean squared errors (MSE) in the magnetic correlations of Hydrogen chains Left: The MSE spin-spin correlation errors for an eight atom hydrogen chain (1.51.5Å  symmetric stretching) between the first atom and its $i^{\text{th}}$ nearest neighbor in a STO-3G basis. Ten sets of 40,960 shots are used to calculate the MSE from the exact statevector, comparing a direct sampling of qubitwise grouped Pauli strings (purple) and a mapping of the shadow state to a CCSD ansatz (orange). Right: The RMS error in the local (same-atom) spin-spin correlation function sampled over 40,960 shots for an Natom hydrogen chain of increasing length.
To further characterize the errors from the property estimates of Sec. III, we can consider the mean squared errors (MSE) in the magnetic correlation functions of Fig. 3 in the main text. This allows us to combine the bias and variance contribution to the errors of the shadow state mapped to the coupled cluster form, and compare directly to the corresponding MSE of the Pauli estimator of these properties for a given shot budget. In Fig. 7, we show these MSE errors in the two-point magnetic correlations, both between increasingly distant atom pairs for a fixed chain length (left plot), and for the same-atom magnetic correlations but with an increasing length of the overall chain. Both estimators sampled the same number of shots (40,960) from the exact ground statevector.
We find that the MSE errors in these magnetic correlations increase approximately linearly with the distance between the atomic sites considered, with RMS errors equivalent between the two approaches from a distance of approximately four sites. This is largely driven from the increasing bias in the CCSD ansatz, also observed in Fig. 3. However, this bias is minimal for same-spin or close neighbour correlation functions, where significant advantages over the Pauli sampling are found. This advantage is only compounded for increasing system sizes (right plot), where we find an RMS error which is largely independent of system size, in contrast to the rapidly increasing MSE error of the Pauli expectation (driven by the increasing variance of this estimator).
Appendix D Calculation of stabilizer overlaps
@@
[…]
@@
Appendix E Direct configurational sampling
Figure 8: Variance scaling for mixed and Pauli estimators of Hydrogen chains over changing correlation strengths and chain length, analogous to Fig. 1, but including a direct sampling of excitation amplitudes in the mixed estimator. (a) Comparison of the distributions of the Pauli and mixed energy estimators, comparing classical shadows to the direct evaluation of the excitation amplitudes via diagonalization and measurement of the operators in Eq. 32 (green), for the ten atom STO-3G basis Hydrogen chain across a range of interatomic distances. The width of the error bar is the standard deviation of the distribution of energies from $1000$ measurements of the exact statevector. (b) The dependence of the standard deviation of each energy estimator with system size, for a fixed atomic distance of 1.51.5Åand $1000$ shots. The black line again shows the Pauli operators grouped into as few commuting sets as possible, while purple is the same for grouping Pauli operators based on qubit wise commutivity. The orange and green lines represent the standards deviations of the mixed estimator using the direct and Clifford shadows methods of calculating the $c_{ij}^{ab}$ amplitudes.